OpenStudy (anonymous):

Evaluate the polynomial 3x^2+11xy+y^2 for x=4 and y=3

OpenStudy (anonymous):

Substitute the value of x and y directly. 3(4^2)+11(4)(3) + (3^2)=? 3(16) + 11(12) + 9=? 48 + 132 +9=? You do the remaining equation. :)

OpenStudy (anonymous):

Complete so as I can learn from it ,Am new to it

OpenStudy (anonymous):

Okay so 48 + 132 +9=? 180 + 9 =? 189 is the answer. The technique is to simply the exponent first and then do the remaining operations. :)

OpenStudy (thejax):